# Binarization-Aware Adjuster: A Theoretical Framework for Bridging Continuous Optimization and Discrete Inference with Application to Edge Detection

### v2

@@ 3.2 Adjusted Loss Function Formulation
    If an element is correctly classified, then the loss on it is 0.
@@
    If an element is incorrectly classified, then the loss on it is positive.
This adjuster emphasizes only misclassified elements. However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization. Therefore, a smoother, differentiable formulation is necessary.
3.3 Design of the Binarization-Aware Adjuster
Let thrthr denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
  • 1.
    C1(Continuity)C1\ (Continuity): The adjuster $Adj$ must be continuous to allow gradient-based optimization.
  • 2.
    C2(Nonnegativity)C2\ (Non-negativity), $Adj(Pred_{i},Gt_{i})$ should be non-negative since it represents the weight of an element. Without loss of generality, assume that $Adj(Pred_{i},Gt_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
@@ 3.3 Design of the Binarization-Aware Adjuster
    C3(UncertaintyEmphasis)C3\ (Uncertainty\ Emphasis): If an element is correctly classified, then the closer the prediction is to thrthr, the higher the weight $Adj(Pred_{i},Gt_{i})$ of it should be, reflecting the loss caused by the uncertainty.
@@
    C4(ConfidenceSaturation)C4\ (Confidence\ Saturation): If a n element is correctly classified with high confidence (i.e., prediction far from thrthr), then $Adj(Pred_{i},Gt_{i})$ should decay to 0 quickly as $Pred_{i}$ goes away from thrthr, representing that further optimizing provide little benefits. Indeed, $Adj(Pred_{i},Gt_{i})$ should close to the maximal value when $Pred_{i}$ is close to thrthr, remain large when $Pred_{i}$ is not far from thrthr, and decay to 0 after some pre-defined distance thr_devthr\_dev.
@@
    C5(ClassificationSensitivity)C5\ (Classification\ Sensitivity): If an element is misclassified, for example if it is an edge one but $Pred_{i}<thr$ or is a non-edge one but $Pred_{i}\geq thr$ in the ED task, its weight $Adj(Pred_{i},Gt_{i})$ should be maximal, representing correctly classification is of the first priority.
These conditions ensure that misclassified elements always receive full attention, and optimization of correct elements is focused on the uncertain parts while conserving effort on highly confident correct parts. We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short. Namely:
@@
A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued at [0,1] satisfying the above conditions C1 to C5.
Let thr_devthr\_dev denote the threshold window from $C4$. To construct a BAA, we define:
@@
A function $f_{thr\_dev}(x)$ is said to be a Distance Weight Function (DWF) regarding the threshold window $thr\_dev\in\mathbb{R}_{+}$, if it satisfies the following properties:
  • 1.
    Continuity: $f_{thr\_dev}(x):[0,thr\_dev]\longrightarrow[0,1]$ is continuous.
  • 2.
    Monotonically Decreasing: $f_{thr\_dev}(x_{1})\leq f_{thr\_dev}(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ with $x_{1}\geq x_{2}$. (or equivalently, $\frac{df_{thr\_dev}}{dx}(x)\leq 0$ for differentiable $f_{the\_dev}(x)$).
  • 3.
    Concavity: $f(\lambda x_{1}+(1-\lambda)x_{2})\geq\lambda f(x_{1})+(1-\lambda)f(x_{2})$, $\forall x_{1},x_{2}\in[0,thr\_dev]$ and $\lambda\in[0,1]$. (or equivalently, $\frac{d^{2}f_{thr\_dev}}{dx^{2}}(x)\leq 0$ for secondary differentiable $f_{the\_dev}(x)$).
  • 4.
    Boundary Conditions: $f_{thr\_dev}(0)=1$, and $f_{thr\_dev}(thr\_dev)=0$.
Then:
$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|Pred_{i}-thr|)&0\leq|Pred_{i}-thr|\leq thr\_dev\\ &0\ &|Pred_{i}-thr|>thr\_dev\end{matrix}\right$. (5)
satisfies C1C4C1-C4.
An example of the DWF $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ can be constructed as:
$f_{b,thr\_dev}(x)=\frac{e^{b\times x}-e^{b\times thr\_dev}}{1-e^{b\times thr\_dev}}$ (6)
where $b\geq 0$ is pre-defined, controlling the decay rate.
To also meet $C5$, we define the masked distance as:
@@
Now, a BAA can be defined as:
$BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=\hat{f}_{thr\_dev}(MD_{thr}(Pred_{i},Gt_{i}))$ (8)
@@
$\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}&1&x<0\\ &f_{thr\_dev}(x)&0\leq x\leq thr\_dev\\ &0&x>thr\_dev\end{matrix}\right$. (9)
@@ 3.4 Theoretical Properties of the BAA
In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 6, generalizes the baseline loss $L$. It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided.. These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation.
@@ 4 Experiment: Application to Edge Detection
To validate the effectiveness of the proposed BAA framework, we apply it to the ED task, which is selected since it is one of the most fundamental tasks in computer vision, supporting substantial high-level tasks, and many tasks can be reformulated to it.
@@ 4.1 Experimental Setting
We adopt the adjuster designed by Equation 6, 7, 8, and 9 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
@@ 4.1.1 Baseline and Benchmark
The proposed method is evaluated across a range of ED models, including: HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], the EES framework integrating the above three models with the double-size selector. Experiments are conducted on four benchmark datasets: BRIND [6], BSDS500 [14], UDED [19], and BIPED2 [18].
All models are retrained from scratch on each dataset to ensure fairness. For consistency and simplicity, only the final output of each model is used for supervision, bypassing any need for hyperparameter tuning in intermediate layers. Each model is trained under the following three configurations of loss function for comparison:
@@ 4.1.2 Dataset Partitions and Augmentation
Dataset preprocessing and postprocessing procedures follow those introduced in [10]. Details are stated as follows.
@@ 4.2 Experimental Results
The results clearly demonstrate that incorporating the proposed BAA method, whether with the fixed threshold (BAA-0.7) or the self-adaptive strategy (BAA-SA), leads to improvements in performance across most models and datasets.
Both variants of the BAA loss improve the baseline WBCE loss across all models and datasets. The self-adaptive version (BAA-SA) tends to achieve slightly better performance than the fixed threshold in most cases.
@@
These results strongly validate the effectiveness of the proposed BAA method, especially the self-adaptive variant. The method is robust across different backbone architectures and datasets, offering consistent improvements with minimal modifications to the existing training pipeline.
@@ 4.3 Ablation Study
This ablation study shows that across all tested configurations, thr_dev=0.2thr\_dev=0.2 consistently yields the best performance. Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
Figure 2: ODS and OIS across thr_devthr\_dev: Figures show how ODS and OIS vary with thr_devthr\_dev values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32). Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss. Threshold thrthr is fixed at 0.7 and $\delta$ at 1.
@@ 5 Discussion
Although the BAA is initially designed for binary decision tasks, its formulation can be naturally extended to general multi-valued discrete inference. A straightforward strategy is to apply the adjuster to each decision boundary separately. That is, whenever a prediction approaches a discrete decision threshold, regardless of the number of categories, a higher optimization weight is assigned. The underlying theoretical principle remains consistent with the binary case established in Section 3, as each boundary can be locally treated as a binarization problem.
While Section 4 presents experimental validation using ED as an example task, the proposed BAA is fundamentally applicable to a broader range of discrete prediction tasks. For instance, experimental results on the BSDS500 dataset, which is also used for contour detection and segmentation, implicitly demonstrate its broader potential. However, due to the substantial computational cost required for large-scale experiments [^3] The experiments on ED alone took approximately three months with GPUs. and the fact that this work primarily aims to establish the theoretical framework rather than exhaustive empirical study, we leave extensive cross-task evaluation for future research. Nevertheless, given that many structured prediction tasks in computer vision, such as segmentation, object detection, and boundary extraction, can be reformulated in terms of ED, we anticipate that the proposed method holds significant potential beyond.
A limitation of the current self-adjusted BAA framework is that model training must be performed twice to determine the binarization threshold thrthr, incurring additional computational overhead during optimization, though it does not affect inference efficiency. Furthermore, the optimal decision threshold may vary across dataset partitions or benchmarks, requiring recalibration when transferring the model. This issue, however, is fundamentally induced by dataset-dependent evaluation settings rather than by the BAA framework itself. The self-adaptive threshold selection still provides a practical mechanism for choosing a reasonable threshold.
Another drawback is that the self-estimated threshold may not precisely match the true optimal inference threshold. For example, in the ED task, the ideal threshold based on the ODS metric is obtained by test data, while the pre-trained threshold might be derived from it. Since the optimal threshold is inherently impossible to determine during training, the self-adaptive estimation should be regarded as a practically effective surrogate rather than an exact optimal value. Nevertheless, such a surrogate can also provide a method to estimate the real optimal decision threshold, since the real one is not practically accessible.
In summary, although additional calibration is required and further task-specific validation remains open, the theoretical framework provides a principled foundation for improving optimization–inference alignment across a wide spectrum of discrete decision problems.
@@ 6 Conclusion
This paper tackles a fundamental yet underexplored issue occurring in discrete decision machine learning tasks: the misalignment between the continuous-valued outputs optimized during training and the discrete predictions required at inference and evaluation. This discrepancy, often overlooked due to the discontinuity of discretization, undermines the alignment between training objectives and final task performance.
To address this, we propose a principled binarization adjustment mechanism that incorporates binarization awareness into the training process. Guided by the Distance Weight Function (DWF), the Binarization-Aware Adjuster (BAA) adaptively modulates each element’s contribution to the loss based on its prediction correctness and confidence. It downweights confidently correct predictions since further optimizing them has little effect on the binary output, while emphasizing ambiguous or misclassified regions that are more critical to the final binarization decision. The resulting adjusted loss function from the BAA generalizes traditional loss formulations and embeds binarization behavior directly into the optimization process. In addition, we introduce a self-adaptive training procedure to dynamically estimate an appropriate binarization threshold for the BAA, further aligning training dynamics with inference behavior. Our method is applied to the edge detection (ED) task, and extensive experiments across representative models and datasets validate its effectiveness and robustness.
In summary, the BAA introduces a theoretically grounded mechanism for bridging continuous optimization and discrete evaluation, with application to the ED task, demonstrating the effectiveness experimentally. And beyond ED task and binarization decision tasks, the principle can inform related discrete structured-prediction tasks, offering a promising direction for broader applications.

### v3

@@ 3.3 Design of the Binarization-Aware Adjuster
    C1 (Continuity): $Adj$ must be continuous to support gradient-based optimization.
@@
    C2 (Non-negativity): $Adj(Pred_{i},Gt_{i})\in[0,1]$, as it represents a weighting factor.
@@
    C3 (Uncertainty Emphasis): For correctly classified elements, the weight should decrease as the prediction goes away from thrthr.
@@
    C4 (Confidence Saturation): For confidently correct elements far from thrthr, the weight should rapidly decay, become zero beyond a predefined distance thr_devthr\_dev.
@@
    C5 (Classification Sensitivity): Misclassified elements should receive the maximal weight.
An adjuster satisfying these properties is referred to as a Binarization-Aware Adjuster (BAA).
@@
A Binarization-Aware Adjuster (BAA) with respect to $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued in $[0,1]$ satisfying conditions C1–C5.
To construct such an adjuster, we introduce the Distance Weight Function (DWF).
@@
A function $f_{thr\_dev}(x)$ is a Distance Weight Function (DWF) with threshold window $thr\_dev\in\mathbb{R}_{+}$ if it satisfies continuity, monotonic decrease, concavity, and boundary conditions: $f_{thr\_dev}(0)=1$ and $f_{thr\_dev}(thr\_dev)=0$.
Using a DWF, an intermediate adjuster satisfying C1–C4 can be defined as:
$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}f_{thr\_dev}(|Pred_{i}-thr|),&|Pred_{i}-thr|\leq thr\_dev,\\ 0,&|Pred_{i}-thr|>thr\_dev.\end{matrix}\right$. (5)
A concrete example of DWF is:
$f_{thr\_dev}(x):=f_{b,thr\_dev}(x):=\frac{e^{b\times x}-e^{b\times thr\_dev}}{1-e^{b\times thr\_dev}},$ (6)
where $b\geq 0$ controls the decay rate.
To additionally satisfy C5, we define the masked distance:
@@
The final BAA is then defined as:
$BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=\hat{f}_{thr\_dev}(MD_{thr}(Pred_{i},Gt_{i})),$ (8)
@@
$\hat{f}_{thr\_dev}(x)=\left\{\begin{matrix}1,&x<0,\\ f_{thr\_dev}(x),&0\leq x\leq thr\_dev,\\ 0,&x>thr\_dev.\end{matrix}\right$. (9)
@@ 3.4 Theoretical Properties of the BAA
In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 6, generalizes the baseline loss $L$. It degenerates to the baseline loss (up to a scale factor) when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains continuity. These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation.
@@ 4 Experiment: Application to Edge Detection
To validate the effectiveness of the proposed BAA framework, we apply it to the ED task. ED is selected as it is a fundamental problem in computer vision, serving as a core component for many higher-level vision tasks and providing a clear setting for evaluating binary decision learning.
@@ 4.1 Experimental Setting
We use weighted binary cross-entropy (WBCE) as the baseline loss function $L$, due to its widespread adoption in ED tasks, and adopt the adjuster defined in Equations 6, 7, 8, and 9 as the BAA.
@@ 4.1.1 Baseline and Benchmark
The proposed method is evaluated on a representative set of ED models, including HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], where EES3-Double denotes the EES framework integrating the above three models with a double-size selector.
Experiments are conducted on four benchmark datasets: BRIND [6], BSDS500 [14], UDED [19], and BIPED2 [18].
All models are retrained from scratch on each dataset to ensure fairness. For consistency and simplicity, only the final output of each model is used for supervision, bypassing any need for hyperparameter tuning in intermediate layers. Each model is trained under the following three loss configurations for comparison:
@@ 4.1.2 Dataset Partitions and Augmentation
Dataset preprocessing and postprocessing procedures follow those introduced in [10]. The details are summarized as follows.
@@ 4.2 Experimental Results
The results indicate that incorporating the proposed BAA method, either with a fixed threshold (BAA-0.7) or with the self-adaptive setting (BAA-SA), generally leads to improved performance across most evaluated models and datasets. Both variants of the BAA loss outperform the baseline WBCE loss in the majority of evaluated settings. The self-adaptive variant (BAA-SA) achieves slightly higher scores than the fixed-threshold version in several cases.
@@
Overall, these results support the effectiveness of the proposed BAA method in the context of ED. The performance gains are consistent across different backbone architectures and datasets, while requiring no modification to model architectures or inference procedures.
@@ 4.3 Ablation Study
This ablation study shows that across all tested configurations, thr_dev=0.2thr\_dev=0.2 consistently yields the best performance. Moreover, the proposed BAA method is robust to a wide range of hyperparameter settings, suggesting that it is not overly sensitive to hyperparameter choices in the ED setting and that a unified configuration can be adopted in practice.
Figure 2: ODS and OIS across thr_devthr\_dev: ODS and OIS values under different thr_devthr\_dev settings (0.1 to 0.7) and decay rates $b$ (8, 16, 32). Experiments are conducted on BRIND using the EES3-Double model with the adjusted WBCE loss. The threshold thrthr is fixed at 0.7 and $\delta$ at 1.
@@ 5 Discussion
Although the BAA is formulated for binary decision tasks, its underlying principle is not restricted to binarization itself. Conceptually, the framework can be extended to multi-valued discrete inference by applying the adjuster around individual decision boundaries. In such cases, when a correct prediction goes away from any discrete boundary, a decreased optimization weight can be assigned accordingly. From a theoretical perspective, this extension can even be realized by the binary formulation established in Section 3, as each local decision boundary can be treated as a binary decision problem.
In this work, experimental validation is conducted using ED as a representative task. This choice is motivated by the clear binary inference rule in ED and its suitability for analyzing optimization-inference misalignment. The primary goal of this study is to establish a theoretically grounded adjustment mechanism rather than to provide an exhaustive empirical evaluation across tasks. While ED also underlies several related vision problems, we deliberately limit our experiments to this setting to maintain a focused and controlled analysis. Further empirical investigations on other tasks are therefore left as future work.
One practical consideration of the BAA formulation is that additional model training for the binarization threshold estimation introduces additional training cost, despite not affecting inference efficiency. However, this overhead is not inherent to the BAA framework itself, but rather stems from the fact that the optimal decision threshold is not accessible in practical tasks. From this perspective, the two-stage estimation process should also be regarded as a general-purpose strategy for approximating decision thresholds in discrete inference problems, rather than as a limitation specific to the BAA framework.
Another limitation is that the estimated threshold may not coincide exactly with the true optimal inference threshold. For instance, in ED, the optimal threshold under the ODS metric is determined using test data, which is inherently inaccessible during training. As a result, the self-adaptive threshold should be interpreted as an effective surrogate rather than an exact optimum. Nevertheless, such a surrogate provides a principled and practically feasible approximation for guiding binarization-aware optimization.
Overall, despite the need for threshold calibration and the focus on a single task, the proposed framework establishes a principled foundation for improving the alignment between continuous optimization and discrete inference in decision-based learning problems.
@@ 6 Conclusion
This paper addresses a fundamental issue in discrete decision learning: the misalignment between continuous-valued optimization during training and discrete predictions at inference, which arises from the discontinuity of discretization operations that are excluded from standard learning pipelines. To address this issue, we introduce a theoretically grounded binarization-aware adjustment (BAA) mechanism, guided by the Distance Weight Function (DWF). BAA incorporates binarization characteristics into continuous optimization by modulating loss contributions according to prediction correctness and proximity to the decision threshold, emphasizing decision-critical elements and downweighting those with limited influence. The proposed framework is applied to the edge detection (ED) task, where binary inference is misaligned with the standard model training. Experimental results on representative models and datasets show that adopting BAA into optimization consistently improves performance. Overall, the BAA provides a principled mechanism for bridging continuous optimization and discrete evaluation, offering a theoretically sound direction for improving decision-aligned learning in discrete inference tasks.